4. (12 points ) Solve:
Answer #1

4. Let A=10 l and flt) = 1 1-1 od 1. The characteristic polynomial of A is. A() = 72-(faceA) + delta = y2 +1: 5 4(2)=0 + 1 =fet fint - lost P=1 I first I-lost fint forst] Got Cost 4 - sl-t host t font 2 t fint tlost] 1 -list of + S fint at | Istrist2103 [3 - 17-449-89 = 11 => 1942] -1] => [= ? and.q=4 listy +2 hent- *1 = 12wt + bont -t). - G+27 i-last 1+2 i 2 sint-lost +2